OpenStudy (anonymous):

find the range of y=-cosx+3 when x is between 0 and pi

OpenStudy (solomonzelman):

Cos(0)=1 Cos(180)=1 the rest are fractions smaller than 1.

OpenStudy (solomonzelman):

my bad, cos(180)=-1

OpenStudy (anonymous):

oh okay i got it!

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \bf y=-cos(0)+3 =-(1)+3=2 }\) \(\large\color{black}{ \bf y=-cos(180)+3 =-(-1)+3=4 }\) you won't get cosine that is \(\large\color{black}{ \bf >│1│ }\)

OpenStudy (solomonzelman):

I mean any other cosine.